# Device-independent certification of desirable properties with a confidence interval

### v1

@@ III Methodologies for hypothesis testing
Hypothesis 1.
[…]
@@
Then, we apply appropriate methods for this kind of hypothesis testing on the collected data to determine an upper bound $\mathcal{P}$ on the $p$-value associated with the hypothesis $\mathcal{H}$. Since a $p$-value quantifies the plausibility of observing the given data when $\mathcal{H}$ holds, a small value of $\mathcal{P}$, say, less than $5\%$ provides a strong indication that $\mathcal{H}$ is falsified. It then follows that the desired feature corresponding to the negation of $\mathcal{H}$ is certified with a confidence $\gamma$ of at least $1-\mathcal{P}$.
@@
where $p^{\text{(prot)}}$ is the $p$-value (upper bound) deduced from some protocol (abbreviated as “prot”). From the definition, it is evident that in the i.i.d. setting, a fewer number of trials is required to achieve the same level of statistical confidence if the corresponding $G^{\text{(prot)}}$ is higher. Next, let us elaborate on the two hypothesis-testing protocols considered in this work.
@@ III.1 Martingale-based protocol
$\left\langle I(v)\right\rangle:=\left\langle\tfrac{\beta^{ab}_{xy}}{P_{xy}}\right\rangle=\sum_{a,b,x,y}\beta^{ab}_{xy}P(ab|xy)\overset{\mathcal{H}}{\leq}B_{\mathcal{H}}(\vec{\beta})$ (24)
may be violated by some ideal quantum correlation $\vec{P}=\vec{P}_{\mathcal{Q}}$ [cf. Eq. 6] to be prepared in an experiment.
Let $v_{j}=(a,b,x,y)_{j}$ be the value realized for the random variables of the measurement outcomes and settings at the $j$-th experimental trial, and $I(v_{j})$ the corresponding value of Bell function for that trial. Moreover, let $\mathbf{v}=\{v_{1},...,v_{i},...,v_{N}\}$. Then, from the observed average value of $I(v)$ over $N$ trials, i.e., $\hat{I}(\mathbf{v})=\sum_{j=1}^{N}\frac{I(v_{j})}{N}$, the following $p$-value upper bound is known [89] to hold whenever $\hat{I}\geq B_{\mathcal{H}}$:
$p^{\text{(mart)}}\leq\left[\left(\frac{b_{+}-B_{\mathcal{H}}}{b_{+}-\hat{I}}\right)^{\frac{b_{+}-\hat{I}}{b_{+}-b_{-}}}\left(\frac{B_{\mathcal{H}}-b_{-}}{\hat{I}-b_{-}}\right)^{\frac{\hat{I}-b_{-}}{b_{+}-b_{-}}}\right]^{N},$ (25)
where, for simplicity, we have suppressed the dependency of $B_{\mathcal{H}}$ on $\vec{\beta}$ (and $\hat{I}$ on $\mathbf{v}$), while the minimum and maximum value of $I(v)$ satisfy
$b_{-}:=\inf_{v}I(v)<B_{\mathcal{H}}\leq\hat{I}<b_{+}:=\sup_{v}I(v).$ (26)
@@
$G^{\text{(mart)}}=\frac{b_{+}-I_{\mathcal{Q}}}{u-b_{-}}\log_{2}\frac{b_{+}-I_{\mathcal{Q}}}{b_{+}-B_{\mathcal{H}}}+\frac{I_{\mathcal{Q}}-b_{-}}{b_{+}-b_{-}}\log_{2}\frac{I_{\mathcal{Q}}-b_{-}}{B_{\mathcal{H}}-b_{-}}.$ (27)
@@ III.2 The prediction-based-ratio (PBR) protocol
Then, the key idea of the PBR protocol is to use this relative frequency $\vec{f}=\{f(ab|xy)\}$ to obtain an optimized Bell-like inequality [^3] Here, the inequality is optimized in the sense that it provides the largest possible asymptotic confidence-gain rate, cf. Eq. 23. and apply that to $v_{i}$ from $i=\{N_{\text{est}}+1,N_{\text{est}}+2,\cdots,N_{\text{est}}+N_{\text{test}}\}$. To this end, we minimize the Kullback-Leibler (KL) divergence [90] from a regularized relative frequency $\vec{f}_{\text{reg}}$ (explained below) to the set $\mathcal{S}_{\mathcal{H}}$ of correlations compatible with $\mathcal{H}$:
@@
The unique [35] minimizer $\vec{P}^{\star}\in\mathcal{S}_{\mathcal{H}}$ can then be used to define the non-negative prediction-based-ratio (PBR)
$R(a,b,x,y):=\frac{f_{\text{reg}}(ab|xy)}{P^{\star}(ab|xy)},$ (30)
@@
Next, notice that $\vec{f}^{\prime}$ typically cannot be cast in the form of Eq. 6. Consequently, we observe empirically that the $R$ obtained by solving Eq. 29 with $\vec{f}^{\prime}$ in place of $\vec{f}_{\text{reg}}$ gives evidently suboptimal performance (see, e.g., Figs. 8, 11 and 13 in Appendix A for some explicit examples). As such, we shall first regularize [35] $\vec{f}^{\prime}$ to some outer approximation of the quantum set $\mathcal{Q}_{\ell}$ by solving Eq. 29 with $\mathcal{S}_{\mathcal{H}}$ replaced by $\mathcal{Q}_{\ell}$. In our work, $\mathcal{Q}_{\ell}$ is the $-\ell$ outer approximation of the quantum set $\mathcal{Q}$ introduced in [16]. However, one may also consider other approximations [65, 21]. Again, for the properties discussed in Section II.3, this regularization process is a conic program (see also [35]). The resulting minimizer, which we call the regularized relative frequency $\vec{f}_{\text{reg}}$ is then fed into Eq. 29 to obtain the desired PBR.
Another important feature of the PBR protocol is that the optimized inequality characterized by $\vec{R}=\{R(a,b,x,y)\}$ can be updated as more data is incorporated into the analysis. In principle, one can update $\vec{R}$ as frequently as one desires. However, this is neither necessary nor efficient. As such, we work with blocks of $N_{\text{blk}}$ trials. The first block of data is used exclusively for producing the first regularized relative frequency, the first PBR $\vec{R}_{1}$, and by applying to the second block of $\mathbf{v}$, we get the test statistic
@@
where $N^{(k)}_{\text{test}}=N_{\text{int}}$ for all $k$. In the next iteration, we determine the PBR $\vec{R}_{2}$ by solving Eq. 29 using $\mathbf{v}$ from the first two blocks, and apply this updated PBR to the third block of $\mathbf{v}$ to get, for $k=2$,
@@ IV Device-independent certification with a confidence interval
We are now ready to present our simulations results involving a finite number of trials. Throughout this section, the results presented consist of an average over $30$ complete Bell tests, each involving $N_{\text{tot}}$ trials. Moreover, we always consider a uniform distribution for (possibly a restricted set of) measurement settings. In each Bell test, we then simulate the raw data $\mathbf{v}=\{v_{i}=(a_{i},b_{i},x_{i},y_{i})\}_{i=1}^{N_{\text{tot}}}$ using the function $\mathsf{sample\_hist}$ from the Lightspeed Matlab toolbox [92]. For the certification with finite data, we set a confidence level of $\gamma=0.99$. We also present some related confidence-gain rates in the respective subsections.
@@ IV.1.1 Negativity certification
Our first example consists of a Bell test based on the quantum strategy presented in Section II.2.3, which leads to a CHSH Bell value of $\mathcal{S}_{\text{\tiny CHSH}}=2\sqrt{2}$. Using Eq. 10, we know that the resulting quantum correlation $\vec{P}_{\text{CHSH}}$ gives a tight negativity lower bound of $\frac{1}{2}$ for a Bell state. From the numerically simulated data, we then perform composite hypothesis testing for the null hypothesis of 1 with $\mathcal{N}_{0}\in\{0,0.01,\cdots,0.50\}$.
Specifically, for the martingale-based protocol, we use Eq. 25 with the CHSH Bell expression of Eq. 4. In this case, $b_{\pm}=\pm 4$ for the chosen $P_{xy}$ while it follows from Eqs. 4, 10 and 24 that
$\sum_{a,b,x,y}(-1)^{xy+a+b}4\,P(a,b,x,y)\overset{\mathcal{N}\leq\mathcal{N}_{0}}{\leq}2+4\mathcal{N}_{0}(\sqrt{2}-1).$ (37)
On the other hand, for the PBR protocol, the optimizing distribution $P^{\star}_{k}(a,b|x,y)$ for the $k$-iteration can be obtained by solving [cf. Eq. 29]
@@
where $\operatorname*{\text{argmin}}_{\vec{P}}$ seeks for the argument minimizing the expression in Eq. 38a, $\vec{f}_{\text{reg}}^{\,(k)}$ is the regularized frequency obtained for the same iteration, and each $P(ab|xy)$ also appears as an optimization variable in the moment matrix $\chi_{\ell}[\rho]$. Then, the PBR used in the computation of $t_{k}$ can be evaluated by replacing $f_{\text{reg}}(ab|xy)$ and $P^{\star}(ab|xy)$, respectively, by $f_{\text{reg}}^{(k)}(ab|xy)$ and $P^{\star}_{k}(ab|xy)$ in Eq. 30. In computing these $t_{k}$, we consider a block size of $N_{\text{blk}}=500$.
@@
Figure 2: Average amount of negativity certifiable from the data observed in a Bell test generating $\vec{P}_{\text{CHSH}}$. For the martingale-based protocol and any given $\mathcal{N}_{0}$ among $\bm{N}_{0}=\{0,0.01,\cdots,0.49\}$, we use Eq. 37 in Eq. 25 to upper-bound $p^{\text{(mart)}}$ after every block of $N_{\text{blk}}=500$ trials, thereby generating $200\times 50$ upper bounds on $p^{\text{(mart)}}$ for a complete Bell test. For the PBR protocol and a given $\mathcal{N}_{0}$ from $\bm{N}_{0}$, we solve Section IV.1.1 by considering the same block size and the level-$3$ outer approximation of $\mathcal{Q}$ introduced in [16]. Then, we obtain $199\times 50$ upper bounds on $p^{\text{(pbr)}}$ from Eqs. 34, 35 and 32. To determine the lower bound on the underlying $\mathcal{N}(\rho)$ with the desired confidence of $\gamma\geq 99\%$, we look for the largest $\mathcal{N}_{0}$ in $\bm{N}_{0}$ such that $\mathcal{H}_{\mathcal{N}(\rho)\leq\mathcal{N}_{0}}$ is rejected with a $p$-value bound being less than or equal to $0.01$. Each data point shown in the plot is an average over $30$ such lower bounds, and the error bar (standard deviation) gives an indication of the spread of the certifiable negativity. To avoid cluttering the plots, in each line, we show only a small number of markers.
@@
These results clearly suggest that the CHSH Bell function of Eq. 4 is optimal for certifying the underlying negativity of |$\ket{\psi_{\text{\tiny MES}}}$ using the martingale-based protocol. Indeed, a separate computation of Eq. 29 and Eq. 30 using $\vec{P}_{\text{CHSH}}$ in place of $\vec{f}_{\text{reg}}$ shows that the optimized Bell-like inequality is exactly equivalent to Eq. 10. How would things change if we perform a DI certification of the negativity underlying the partially entangled state |$\ket{\psi(\theta)}$, Eq. 17a? To this end, consider the quantum strategy of Section II.2.3, whose resulting correlation $\vec{P}_{\theta}$ gives the maximal Bell CHSH violation for |$\ket{\psi(\theta)}$, as well as the maximal violation of the tilted CHSH Bell inequality Eq. 19. Then, instead of repeating the same analysis, we show in Fig. 3 the confidence-gain rates due to both protocols for certifying several given fractions of the underlying negativity. From the plots shown, it is evident that asymptotically, the martingale-based protocol employing the CHSH Bell function is far from optimal for certifying the underlying negativity of |$\ket{\psi(\theta)}$. Indeed, the PBR protocol could identify some other Bell-like inequality that gives a much better confidence-gain rate, especially for the correlations arising from |$\ket{\psi(\theta)}$ that is weakly entangled (small $\theta$). To a large extent, this can be understood by noting that the negativity lower bound of Eq. 10 due to its CHSH Bell violation is generally far from tight for these states, see Fig. 10 in Appendix A.
@@
Figure 3: Asymptotic confidence-gain rate $G^{\text{(prot)}}$ based on the family of quantum correlations $\vec{P}_{\theta}$ derived from Section II.2.3, where $\theta=\frac{k\pi}{180}$ rad, $k=\{1,2,\cdots,45\}$, parametrizes the two-qubit entangled state |$\ket{\psi(\theta)}=\cos{\theta}\ket{00}+\sin{\theta}\ket{11}$. Here, we again consider the composite hypothesis 1, with $\mathcal{N}_{0}=\{0.8\mathcal{N}_{\psi(\theta)},0.85\mathcal{N}_{\psi(\theta)},0.9\mathcal{N}_{\psi(\theta)},0.95\mathcal{N}_{\psi(\theta)}\}$ and $\mathcal{N}_{\psi(\theta)}$ being the negativity of $\psi(\theta)$, see Fig. 10. The gain rate for the martingale-based protocol is computed from Eq. 27 using the CHSH Bell-like inequality of Eq. 37 whereas that for the PBR protocol is evaluated from Eq. 36 using the correlation derived from Section II.2.3.

### v2

@@ III Methodologies for hypothesis testing
Null Hypothesis 1.
$\mathcal{H}_{\mathcal{N}(\rho)\leq\mathcal{N}_{0}}$: In every experimental trial, the underlying state has a negativity less than or equal to $\mathcal{N}_{0}$.
@@
Then, we apply appropriate methods for this kind of hypothesis testing on the collected data to determine an upper bound $\mathfrak{p}$ on the $p$-value associated with the hypothesis $\mathcal{H}$. Since a $p$-value quantifies the plausibility of observing the given data when $\mathcal{H}$ holds, a small value of $\mathfrak{p}$, say, less than $5\%$ provides a strong indication that $\mathcal{H}$ is falsified. It then follows that the desired feature corresponding to the negation of $\mathcal{H}$ is certified with a confidence $\gamma$ of at least $1-\mathfrak{p}$.
@@
where $p^{\text{(prot)}}$ is the $p$-value (upper bound) deduced from some protocol (abbreviated as “prot”). From the definition, it is evident that asymptotically, and in the i.i.d. setting, a fewer number of trials is required to achieve the same level of statistical confidence if the corresponding $G^{\text{(prot)}}$ is higher. Next, let us elaborate on the two hypothesis-testing protocols considered in this work.
@@ III.1 Martingale-based protocol
$\left\langle I(v)\right\rangle:=\tfrac{\left\langle\beta^{ab}_{xy}\right\rangle}{P_{xy}}=\sum_{a,b,x,y}\beta^{ab}_{xy}P(ab|xy)\overset{\mathcal{H}}{\leq}B_{\mathcal{H}}(\vec{\beta})$ (24)
may be violated by some quantum correlation $\vec{P}=\vec{P}_{\mathcal{Q}}$ [cf. Eq. 6] to be prepared in an experiment.
Let $v_{j}=(a_{j},b_{j},x_{j},y_{j})$ be the value realized for the random variables of the measurement outcomes and settings at the $j$-th experimental trial, and $I(v_{j})$ the corresponding value of Bell function for that trial. Moreover, let $\mathbf{v}=\{v_{1},...,v_{i},...,v_{N}\}$. Then, from the observed average value of $I(v)$ over $N$ trials, i.e., $\hat{I}(\mathbf{v})=\sum_{j=1}^{N}\frac{I(v_{j})}{N}$, the following $p$-value upper bound is known [91] to hold whenever $\hat{I}\geq B_{\mathcal{H}}$:
$p^{\text{(mart)}}\leq\left[\left(\frac{\mathfrak{b}_{+}-B_{\mathcal{H}}}{\mathfrak{b}_{+}-\hat{I}}\right)^{\frac{\mathfrak{b}_{+}-\hat{I}}{\mathfrak{b}_{+}-\mathfrak{b}_{-}}}\left(\frac{B_{\mathcal{H}}-\mathfrak{b}_{-}}{\hat{I}-\mathfrak{b}_{-}}\right)^{\frac{\hat{I}-\mathfrak{b}_{-}}{\mathfrak{b}_{+}-\mathfrak{b}_{-}}}\right]^{N},$ (25)
where, for simplicity, we have suppressed the dependency of $B_{\mathcal{H}}$ on $\vec{\beta}$ (and $\hat{I}$ on $\mathbf{v}$), while the minimum and maximum value of $I(v)$ over all possible values of $v=(a,b,x,y)$ are
$\mathfrak{b}_{-}:=\inf_{v}I(v)<B_{\mathcal{H}}\leq\hat{I}<\mathfrak{b}_{+}:=\sup_{v}I(v).$ (26)
@@
$G^{\text{(mart)}}=\frac{\mathfrak{b}_{+}-I_{\mathcal{Q}}}{\mathfrak{b}_{+}-\mathfrak{b}_{-}}\log_{2}\frac{\mathfrak{b}_{+}-I_{\mathcal{Q}}}{\mathfrak{b}_{+}-B_{\mathcal{H}}}+\frac{I_{\mathcal{Q}}-\mathfrak{b}_{-}}{\mathfrak{b}_{+}-\mathfrak{b}_{-}}\log_{2}\frac{I_{\mathcal{Q}}-\mathfrak{b}_{-}}{B_{\mathcal{H}}-\mathfrak{b}_{-}}.$ (27)
@@ III.2 The prediction-based-ratio (PBR) protocol
The key idea of the PBR protocol is to use this relative frequency $\vec{f}=\{f(ab|xy)\}$ to obtain an optimized Bell-like inequality [^3] Here, the inequality is optimized in the sense that it provides the largest possible asymptotic confidence-gain rate, cf. Eq. 23. and apply that to $v_{i}$ from $i=\{N_{\text{est}}+1,N_{\text{est}}+2,\cdots,N_{\text{est}}+N_{\text{test}}\}$. To this end, we minimize the Kullback-Leibler (KL) divergence [92] from a regularized relative frequency $\vec{f}_{\text{reg}}$ (explained below) to the set $\mathcal{S}_{\mathcal{H}}$ of correlations compatible with $\mathcal{H}$:
@@
The unique [37] minimizer $\vec{P}_{\star}\in\mathcal{S}_{\mathcal{H}}$ can then be used to define the non-negative prediction-based-ratio (PBR)
$R(a,b,x,y):=\frac{f_{\text{reg}}(ab|xy)}{P_{\star}(ab|xy)},$ (30)
@@
Next, notice that $\vec{f}^{\prime}$ typically cannot be cast in the form of Eq. 6. Consequently, we observe empirically that the $R$ obtained by solving Eq. 29 with $\vec{f}^{\prime}$ in place of $\vec{f}_{\text{reg}}$ gives evidently suboptimal performance (see, e.g., Figs. 10, 13 and 15 in Appendix A for some explicit examples). As such, we shall first regularize [37] $\vec{f}^{\prime}$ to some outer approximation of the quantum set $\mathcal{Q}_{\ell}$ by solving Eq. 29 with $\mathcal{S}_{\mathcal{H}}$ replaced by $\mathcal{Q}_{\ell}$. In our work, $\mathcal{Q}_{\ell}$ is the $-\ell$ outer approximation of the quantum set $\mathcal{Q}$ introduced in [17]. However, one may also consider other approximations [67, 22]. Since all these outer approximations admit an SDP characterization, this regularization process is a conic program (see also [37]). The resulting minimizer, which we call the regularized relative frequency $\vec{f}_{\text{reg}}$ is then fed into Eq. 29 to obtain the desired PBR.
Another important feature of the PBR protocol is that the optimized inequality characterized by $\vec{R}=\{R(a,b,x,y)\}$ can be updated as more data is incorporated into the analysis. In principle, one can update $\vec{R}$ as frequently as one desires. However, this is neither necessary nor efficient. As such, we work with blocks of $N_{\text{blk}}$ trials. The first block of data is used exclusively for producing the first regularized relative frequency, the first PBR $\vec{R}_{1}$, and by applying to the second block of $\mathbf{v}$, we get the first test statistic
@@
where $N^{(k)}_{\text{test}}=N_{\text{blk}}$ for all $k$. In the next iteration, we determine the PBR $\vec{R}_{2}$ by solving Eq. 29 using $\mathbf{v}$ from the first two blocks, and apply this updated PBR to the third block of $\mathbf{v}$ to get, for $k=2$,
@@ IV Device-independent certification with a confidence interval
We are now ready to present our simulations results involving a finite number of trials. Throughout this section, the results presented for finite trials consist of an average over $30$ complete Bell tests, each involving $N_{\text{tot}}=10^{5}$ trials, with the trials partitioned into blocks of size $N_{\text{blk}}=500$. Moreover, we always consider a uniform distribution for (possibly a restricted set of) measurement settings. In each Bell test, we then simulate the raw data $\mathbf{v}=\{v_{i}=(a_{i},b_{i},x_{i},y_{i})\}_{i=1}^{N_{\text{tot}}}$ using the function $\mathsf{sample\_hist}$ from the Lightspeed Matlab toolbox [94]. For the certification with finite data, we set a confidence level of $\gamma=0.99$. We also present some related confidence-gain rates in the respective subsections.
@@ IV.1.1 Negativity certification
Our first example consists of a Bell test based on the quantum strategy presented in Eq. 15, which leads to a CHSH Bell value of $\mathcal{S}_{\text{\tiny CHSH}}=2\sqrt{2}$. Using Eq. 10, we know that the resulting quantum correlation $\vec{P}_{\text{CHSH}}$ gives a tight negativity lower bound of $\frac{1}{2}$ for a Bell state. From the numerically simulated data, we then perform composite hypothesis testing for Null Hypothesis 1 with $\mathcal{N}_{0}\in\{0,0.01,\cdots,0.50\}$.
Specifically, for the martingale-based protocol, we use Eq. 25 with the CHSH Bell expression of Eq. 4. In this case, $\mathfrak{b}_{\pm}=\pm 4$ for the chosen $P_{xy}$ while it follows from Eqs. 4, 10 and 24 that
$\sum_{a,b,x,y}(-1)^{xy+a+b}4\,P(abxy)\overset{\mathcal{N}\leq\mathcal{N}_{0}}{\leq}2+4\mathcal{N}_{0}(\sqrt{2}-1).$ (37)
On the other hand, for the PBR protocol, the optimizing distribution $P^{(k)}_{\star}(a,b|x,y)$ for the $k$-iteration can be obtained by solving [cf. Eq. 29]
@@
where $\operatorname*{\text{argmin}}_{\vec{P}}$ seeks for the argument minimizing the expression in Eq. 38a, $\vec{f}_{\text{reg}}^{\,(k)}$ is the regularized frequency obtained for the same iteration, and each $P(ab|xy)$ also appears as an optimization variable in the moment matrix $\chi_{\ell}[\rho]$. Then, the PBR used in the computation of $t_{k}$ can be evaluated by replacing $f_{\text{reg}}(ab|xy)$ and $P_{\star}(ab|xy)$ in Eq. 30, respectively, by $f_{\text{reg}}^{(k)}(ab|xy)$ and $P^{(k)}_{\star}(ab|xy)$.
@@
Figure 2: Negativity certifiable from the data observed in a Bell test generating $\vec{P}_{\text{CHSH}}$. For the martingale-based protocol and any given $\mathcal{N}_{0}$ among $\bm{N}_{0}=\{0,0.01,\cdots,0.49\}$, we use Eq. 37 in Eq. 25 to upper-bound $p^{\text{(mart)}}$ after every block of $N_{\text{blk}}=500$ trials, thereby generating $200\times 50$ upper bounds on $p^{\text{(mart)}}$ for a complete Bell test. For the PBR protocol and a given $\mathcal{N}_{0}$ from $\bm{N}_{0}$, we solve Eq. 38 by considering the same block size and the level-$3$ outer approximation of $\mathcal{Q}$ introduced in [17]. Then, we obtain $199\times 50$ upper bounds on $p^{\text{(pbr)}}$ from Eqs. 34, 35 and 32. To determine the lower bound on the underlying $\mathcal{N}(\rho)$ with the desired confidence of $\gamma\geq 99\%$, we look for the largest $\mathcal{N}_{0}$ in $\bm{N}_{0}$ such that $\mathcal{H}_{\mathcal{N}(\rho)\leq\mathcal{N}_{0}}$ is rejected with a $p$-value bound being less than or equal to $0.01$. Each data point shown in the plot is an average over $30$ such lower bounds, and the error bar (standard deviation) gives an indication of the spread of the certifiable negativity. To avoid cluttering the plots, in each line, we show only a small number of markers.
@@
These results clearly suggest that the CHSH Bell function of Eq. 4 is optimal for certifying the underlying negativity of |$\ket{\psi_{\text{\tiny MES}}}$ using the martingale-based protocol. Indeed, a separate computation of Eq. 29 and Eq. 30 using $\vec{P}_{\text{CHSH}}$ in place of $\vec{f}_{\text{reg}}$ show that, within a precision of $10^{-4}$, the optimized Bell-like inequality for $\mathcal{N}_{0}=0,0.05,\cdots,1$ is equivalent to Eq. 10. How would things change if we perform a DI negativity certification using the data generated from the partially entangled state |$\ket{\psi(\theta)}$, Eq. 17a? To this end, consider the quantum strategy of Eq. 17, whose resulting correlation $\vec{P}_{\theta}$ gives the maximal Bell CHSH violation for |$\ket{\psi(\theta)}$, as well as the maximal violation of the tilted CHSH Bell inequality of Eq. 19. Then, instead of repeating the same analysis, we show in Fig. 3 the confidence-gain rates due to both protocols for certifying several given fractions of the underlying negativity. From the plots shown, it is evident that asymptotically, the martingale-based protocol employing the CHSH Bell function is far from optimal for certifying the underlying negativity of |$\ket{\psi(\theta)}$. Indeed, the PBR protocol could identify some other Bell-like inequality that gives a much better confidence-gain rate, especially for the correlations arising from |$\ket{\psi(\theta)}$ that is weakly entangled (small $\theta$). To a large extent, this can be understood by noting that the negativity lower bound of Eq. 10 due to its CHSH Bell violation is generally far from tight for these states, see Fig. 12 in Appendix A.
@@
Figure 3: Asymptotic confidence-gain rate $G^{\text{(prot)}}$ based on the family of quantum correlations $\vec{P}_{\theta}$ derived from Eq. 17, where $\theta=\frac{k\pi}{180}$ rad, $k=\{1,2,\cdots,45\}$, parametrizes the two-qubit entangled state |$\ket{\psi(\theta)}=\cos{\theta}\ket{00}+\sin{\theta}\ket{11}$. Here, we again consider the composite hypothesis 1, with $\mathcal{N}_{0}=\{0.8\mathcal{N}_{\psi(\theta)},0.85\mathcal{N}_{\psi(\theta)},0.9\mathcal{N}_{\psi(\theta)},0.95\mathcal{N}_{\psi(\theta)}\}$ and $\mathcal{N}_{\psi(\theta)}$ being the negativity of $\psi(\theta)$, see Fig. 12. The gain rate for the martingale-based protocol is computed from Eq. 27 using the CHSH Bell-like inequality of Eq. 37 whereas that for the PBR protocol is evaluated from Eq. 36 using the correlation derived from Eq. 17.